# Cryptocurrency as an Investable Asset Class: Coming of Age

### v3

@@ 3.6 Stylized fact 6: When the funding dries up, we finally learn the worth of futures
Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black & Scholes 1973; Merton 1973; Black 1976).
There is widespread usage of derivatives in the cryptocurrency market, albeit with novel designs. Among all derivatives, perpetual futures are the most common derivative contracts in cryptocurrency markets. For example, Schmeling, Schrimpf & Todorov (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample. Perpetuals, inspired by Shiller (1993), have no maturity; positions are marked to market (typically every 8 hours) and a funding rate—positive or negative by design—flows between longs and shorts to keep futures prices anchored to spot. Ackerer, Hugonnier & Jermann (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures. The crypto-carry trade strategy of Schmeling, Schrimpf & Todorov (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
@@
The figure plots the cumulative returns from Tether Carry Trades (left axis) and from a long buy-and-hold strategy in Bitcoin (right axis). The Carry Trade is defined as in Schmeling, Schrimpf & Todorov (2023). The sample goes from August 1, 2020 to May 31, 2025.
@@
A fundamental insight of information economics is that markets with asymmetric information can fail to allocate resources efficiently. When one party possesses superior information, such as inside information, prices may deviate from fundamentals, and capital may be misallocated (e.g., Akerlof 1970). Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983; Dye 1985; Healy & Palepu 2001; Verrecchia 2001). However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001; Ma & Shleifer 2025). Policymakers thus issue regulations to address these inefficiencies and to produce credible, comparable, and decision-useful information. Intermediaries, such as analysts and auditors, emerge to reduce information asymmetry by analyzing and validating information (Watts & Zimmerman 1983).
@@
On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George & Patatoukas 2021; Liu, Sheng & Wang 2025; Du, Huang & Wu 2025). Howell, Niessner & Yermack (2020) and Bourveau et al. (2022) show the importance of voluntary disclosures in the initial coin offering stage. Studies also examine information intermediaries (e.g., analysts and auditors) and find that they play an important role in this market (e.g., Bourveau et al. 2022; Bourveau, Brendel & Schoenfeld 2024).
@@ 4 ADDITIONAL DISCUSSIONS
Recent research increasingly examines individual investor beliefs and behavior in cryptocurrency markets and compares and contrast them with those in traditional investments. Kogan et al. (2024) show that the trading behaviors of retail investors are drastically different between cryptocurrencies and traditional assets, such as stocks and gold. In particular, retail investors are contrarian in stocks and gold, but trend-chasing in cryptocurrencies. Kogan et al. (2024) conjecture that the trend-chasing behavior is due to cryptocurrency price changes affecting the likelihood of future widespread adoption, which leads retail investors to update their price expectations. Chan et al. (2020) use data from a medium-sized cryptocurrency exchange and document that investor portfolios are typically small in value, highly concentrated in a few coins, and characterized by short holding periods. Weber et al. (2023) complement this evidence using repeated large-scale surveys of U.S. households, showing that cryptocurrency investors differ systematically from other investors. Alvarez, Argente & Van Patten (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption. Han, Liu & Sui (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets. Finally, Benetton & Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies. They find that individual beliefs of cryptocurrency are different and are critical in explaining cryptocurrency valuation.
@@
The bulk of this review is about fungible tokens. A significant segment of the cryptocurrency market, however, is non-fungible tokens (NFTs). Goetzmann & Nozari (2022) are the first to construct repeated sales return indices of NFTs and examine their properties. Borri et al. (2022a) provide a comprehensive analysis of the NFT market. They highlight the importance of visual features in cryptocurrency valuation and how it leads to a conspicuous consumption effect in the market. Furthermore, Borri et al. (2022a) show that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market. Huang & Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing. Oh, Rosen & Zhang (2025) document preference-driven herding in NFT primary market sales.
@@
A stream of work examines the so-called dark side of cryptocurrencies, especially in the early years. Foley, Karlsen & Putniņš (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov & Schoar (2021) that show the results of Foley, Karlsen & Putniņš (2019) are largely driven by double-counting. Griffin & Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation. Amiram, Jørgensen & Rabetti (2022) show that abnormal spikes in on‑chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, Cong et al. (2025) provide an anatomy of crypto-enabled crimes and identify ransomware as the dominant activity, while Cong et al. (2023a) review crypto-related scams, including Ponzi schemes and “rug pulls” in decentralized projects and token offerings. Li, Shin & Wang (2025) document pervasive cryptocurrency pump‑and‑dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre‑event run‑ups consistent with insider gains at outsiders’ expense.

### v4

@@ 3.6 Stylized fact 6: When the funding dries up, we finally learn the worth of futures
Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black and Scholes 1973; Merton 1973; Black 1976).
There is widespread usage of derivatives in the cryptocurrency market, albeit with novel designs. Among all derivatives, perpetual futures are the most common derivative contracts in cryptocurrency markets. For example, Schmeling, Schrimpf, and Todorov (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample. Perpetuals, inspired by Shiller (1993), have no maturity; positions are marked to market (typically every 8 hours) and a funding rate—positive or negative by design—flows between longs and shorts to keep futures prices anchored to spot. Ackerer, Hugonnier, and Jermann (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures. The crypto-carry trade strategy of Schmeling, Schrimpf, and Todorov (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
@@
The figure plots the cumulative returns from Tether Carry Trades (left axis) and from a long buy-and-hold strategy in Bitcoin (right axis). The Carry Trade is defined as in Schmeling, Schrimpf, and Todorov (2023). The sample goes from August 1, 2020 to May 31, 2025.
@@
A fundamental insight of information economics is that markets with asymmetric information can fail to allocate resources efficiently. When one party possesses superior information, such as inside information, prices may deviate from fundamentals, and capital may be misallocated (e.g., Akerlof 1970). Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983; Dye 1985; Healy and Palepu 2001; Verrecchia 2001). However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001; Ma and Shleifer 2025). Policymakers thus issue regulations to address these inefficiencies and to produce credible, comparable, and decision-useful information. Intermediaries, such as analysts and auditors, emerge to reduce information asymmetry by analyzing and validating information (Watts and Zimmerman 1983).
@@
On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George and Patatoukas 2021; Liu, Sheng, and Wang 2025; Du, Huang, and Wu 2025). Howell, Niessner, and Yermack (2020) and Bourveau et al. (2022) show the importance of voluntary disclosures in the initial coin offering stage. Studies also examine information intermediaries (e.g., analysts and auditors) and find that they play an important role in this market (e.g., Bourveau et al. 2022; Bourveau, Brendel, and Schoenfeld 2024).
@@ 4 ADDITIONAL DISCUSSIONS
Recent research increasingly examines individual investor beliefs and behavior in cryptocurrency markets and compares and contrast them with those in traditional investments. Kogan et al. (2024) show that the trading behaviors of retail investors are drastically different between cryptocurrencies and traditional assets, such as stocks and gold. In particular, retail investors are contrarian in stocks and gold, but trend-chasing in cryptocurrencies. Kogan et al. (2024) conjecture that the trend-chasing behavior is due to cryptocurrency price changes affecting the likelihood of future widespread adoption, which leads retail investors to update their price expectations. Chan et al. (2020) use data from a medium-sized cryptocurrency exchange and document that investor portfolios are typically small in value, highly concentrated in a few coins, and characterized by short holding periods. Weber et al. (2023) complement this evidence using repeated large-scale surveys of U.S. households, showing that cryptocurrency investors differ systematically from other investors. Alvarez, Argente, and Van Patten (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption. Han, Liu, and Sui (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets. Finally, Benetton and Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies. They find that individual beliefs of cryptocurrency are different and are critical in explaining cryptocurrency valuation.
@@
The bulk of this review is about fungible tokens. A significant segment of the cryptocurrency market, however, is non-fungible tokens (NFTs). Goetzmann and Nozari (2022) are the first to construct repeated sales return indices of NFTs and examine their properties. Borri et al. (2022a) provide a comprehensive analysis of the NFT market. They highlight the importance of visual features in cryptocurrency valuation and how it leads to a conspicuous consumption effect in the market. Furthermore, Borri et al. (2022a) show that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market. Huang and Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing. Oh, Rosen, and Zhang (2025) document preference-driven herding in NFT primary market sales.
@@
A stream of work examines the so-called dark side of cryptocurrencies, especially in the early years. Foley, Karlsen, and Putniņš (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov and Schoar (2021) that show the results of Foley, Karlsen, and Putniņš (2019) are largely driven by double-counting. Griffin and Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation. Amiram, Jørgensen, and Rabetti (2022) show that abnormal spikes in on‑chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, Cong et al. (2025) provide an anatomy of crypto-enabled crimes and identify ransomware as the dominant activity, while Cong et al. (2023a) review crypto-related scams, including Ponzi schemes and “rug pulls” in decentralized projects and token offerings. Li, Shin, and Wang (2025) document pervasive cryptocurrency pump‑and‑dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre‑event run‑ups consistent with insider gains at outsiders’ expense.